OpenStudy (anonymous):

Solve S = 2LW + 2LH + 2WH for H.

OpenStudy (anonymous):

The first step would be to try to isolate "H" on one side of the equation.

OpenStudy (anonymous):

Or any terms with "H" in them.

OpenStudy (anonymous):

How would I start off?

OpenStudy (anonymous):

S = 2LW + 2LH + 2WH I see that "2LW" does not have an "H" in it, so I will subtract it from both sides. S - 2LW = 2LW - 2LW + 2LH + 2WH S - 2LW = 2LH + 2WH Can you continue?

OpenStudy (anonymous):

Would the answer be S - 2LW / 2LW = H ??

OpenStudy (anonymous):

Hm, I think you're missing a term there. Where is the 2W from 2WH?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

You're missing a term. Can I see your work?

OpenStudy (anonymous):

S = 2LW + 2LH + 2WH S - 2LW = 2LW - 2LW + 2LH + 2WH S - 2LW = 2LH + 2WH S - 2LW / 2LW = H

OpenStudy (anonymous):

There's a mistake from step 3 to 4. Looking at step 3, to isolate "H" what can you do? Or what did you do?

OpenStudy (anonymous):

I divided 2L and 2W to get H alone.

OpenStudy (anonymous):

Have you learned common factoring? You can common factor the "H" out on the right side in step 3. Try that.

OpenStudy (anonymous):

Since both terms have an "H" in them, you can free the "H" by common factoring.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

S = 2LW + 2LH + 2WH S - 2LW = 2LW - 2LW + 2LH + 2WH S - 2LW = 2LH + 2WH S - 2LW = H(2L + 2W) Does this step make sense?

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Lol thank you for being patient with me.. I'm having a little trouble because I've had a weird day. :)

OpenStudy (anonymous):

Don't worry about it, you should see how hopeless I am on the physics section. :P People must pull their hair out while trying to help me.

OpenStudy (anonymous):

Lol. Yikes! Good luck with that!! :D